OpenStudy (anonymous):

what are all the roots of the following polynomial equations x^4+3x^3-5x^2-12x+4=0

OpenStudy (anonymous):

(x-2)(x^3+5x^2+5x-2)=0

OpenStudy (anonymous):

x=2 (x^3+5x^2+5x-2)=0________________/(x+2)(x^2+3x-1)=0 x=-2 or (x^2+3x-1)=0 \[x=\frac{ -3 \pm \sqrt{13}}{ 2}\]